OpenStudy (anonymous):

Solve the differential equation y˝ = 6t + sin t, given y´(0) = 0 and y(0) = 5.

OpenStudy (mathmale):

Solving this d. e. is pretty straightforward, given that you need only integrate twice with respect to t and to evaluate the constants of integration based on the initial conditions.

OpenStudy (mathmale):

In your shoes, I'd write y˝ = 6t + sin t as \[\frac{ d ^{2}y }{ dt ^{2} }= 6t +\sin t\]

OpenStudy (mathmale):

which in turn could be written as \[\frac{ d }{ dt }(\frac{ dy }{ dt }) = 6t + \sin t\]

OpenStudy (mathmale):

Multiply every term here by dt, and then integrate the resulting equation with respect to t. Mind trying that?